XYZ Co. operates in a competitive market. Its marginal product of labor is 1/L, and it takes the

wage and price as given. Derive the firm's short-run demand for labor as a function of w and p.

How much labor will the firm hire if W=₵20 and P=₵100?


1
Expert's answer
2021-10-13T20:43:11-0400

(i)

Assuming:

Price of product =p.

Wage = w.

The input demand for labor is derived as:

"L=f (w,p)"

(ii)

"MP_L=\frac{W}{P}"


"\frac {1}{L}=\frac{W}{P}"


"\frac{1}{L}=\frac{20}{100}"


"0.2L=1"

"L=5."